# Optimistic Query Routing in Clustering-based Approximate Maximum Inner Product Search

### v1

@@ Appendix C Sketching the covariance matrix
Let $\Sigma=D+R$ be the decomposition of the PSD covariance matrix $\Sigma$ into its diagonal $D$ and residual $R=\Sigma-D$. Let $Q\Lambda Q^{\top}$ be the orthogonal eigendecomposition of $D^{-1/2}RD^{-1/2}$. Recall that we define $\tilde{\Sigma}$, the sketch, for some $1\leq t\leq d$ as
$\tilde{\Sigma}:=D+D^{{1}/{2}}[Q]_{t}[\Lambda]_{t}[Q]_{t}^{\top}D^{{1}/{2}}$
We start by providing a proof for Lemma 2 given the assumptions in the lemma are true; i.e. the $(t+1)$-th eigenvalue of $D^{-1/2}\Sigma D^{-1/2}$ is greater than or equal to $2$ and $\min_{i\in[d]}\Sigma_{ii}/\max_{j\in[d]}\Sigma_{jj}\geq(1-\epsilon)$ for some $10\leq\epsilon<$ 1. We then proceed to justify these assumptions in the following section.
C.1 Proof of Lemma 2
@@
First, we state three facts that we will use to simplify our proof for Lemma 2.
@@
Assuming the eigenvalues $\Lambda$ of $D^{-1/2}RD^{-1/2}$ are sorted in non-increasing order, we have that $I+D^{-1/2}RD^{-1/2}=Q(I+\Lambda)Q^{\top}$. In words, the eigenvectors of $I+D^{-1/2}RD^{-1/2}$ are the same as $D^{-1/2}RD^{-1/2}$ and the $i$-th eigenvalue is $\Lambda_{i}+1$.
As a corollary, since $I+D^{-1/2}RD^{-1/2}$ is positive semidefinite, we have that $\Lambda_{i}\geq-1$ for all $i\in[d]$.
@@
Now we are ready to prove the lemma. For any arbitrary vector $v\in\operatorname*{\mathbb{R}}^{d}$ we have that
$v^{\top}(D^{-\frac{1}{2}}\Sigma D^{-\frac{1}{2}}-D^{-\frac{1}{2}}\tilde{\Sigma}D^{-\frac{1}{2}})v$ $=v^{\top}(D^{-\frac{1}{2}}RD^{-\frac{1}{2}}-[Q]_{t}[\Lambda]_{t}[Q]_{t}^{\top})v$ (10)
$=\sum_{i=t+1}^{d}\Lambda_{i}\cdot\langle Q_{i},v\rangle^{2}$ (11)
where the first equality follows by the definition of $\tilde{\Sigma}$ and the second by Fact 1. Hence, by the definition of $\text{err}(\cdot,\cdot)$, we have that $\text{err}(D^{-\frac{1}{2}}\Sigma D^{-\frac{1}{2}},D^{-\frac{1}{2}}\tilde{\Sigma}D^{-\frac{1}{2}})=\max_{l\in[t+1,d]}|\Lambda_{l}|$.
Using Fact 2, we also have that
$\text{err}_{t}(I+D^{-\frac{1}{2}}RD^{-\frac{1}{2}})$ $=\text{err}(I+D^{-\frac{1}{2}}RD^{-\frac{1}{2}},[Q]_{t}[I+\Lambda]_{t}[Q]_{t}^{\top})$
$=\max_{l\in[t+1,d]}|1+\Lambda_{l}|=1+\Lambda_{t+1}\geq 0.$
By our assumption for the lemma, we have that if $1+\Lambda_{t+1}\geq 2$, then $\Lambda_{t+1}\geq 1$. Hence, we have that $\max_{l\in[t+1,d]}|\Lambda_{l}|\leq 1+\Lambda_{t+1}$ since $\Lambda_{l}\geq-1$ for all $l\in[d]$. Hence, when $\Lambda_{t+1}\geq 1$, we have that
$\text{err}(D^{-\frac{1}{2}}\Sigma D^{-\frac{1}{2}},D^{-\frac{1}{2}}\tilde{\Sigma}D^{-\frac{1}{2}})\leq\text{err}_{t}(D^{-\frac{1}{2}}\Sigma D^{-\frac{1}{2}}).$ (12)
In order to prove the lemma, we expand out the definition of $\text{err}(\cdot,\cdot)$
$\text{err}(D^{-\frac{1}{2}}\Sigma D^{-\frac{1}{2}},D^{-\frac{1}{2}}\tilde{\Sigma}D^{-\frac{1}{2}})$ $=\sup_{v\in\operatorname*{\mathbb{R}}^{d}}\frac{|v^{\top}D^{-1/2}(\Sigma-\tilde{\Sigma})D^{-1/2}v|}{\|v\|_{2}^{2}}.$
Since we have that $D$ has strictly positive entries on the diagonal, we can do a change of variables, setting $u=D^{1/2}v$. Denoting $\max_{i\in[d]}D_{ii}$ by $\|D\|_{\infty}$, this gives us
$\text{err}(D^{-\frac{1}{2}}\Sigma D^{-\frac{1}{2}},D^{-\frac{1}{2}}\tilde{\Sigma}D^{-\frac{1}{2}})$ $=\sup_{u\in\operatorname*{\mathbb{R}}^{d}}\frac{|v^{\top}(\Sigma-\tilde{\Sigma})v|}{\|D^{1/2}v\|_{2}^{2}}\geq\frac{1}{\|D\|_{\infty}}\cdot\sup_{u\in\operatorname*{\mathbb{R}}^{d}}\frac{|v^{\top}(\Sigma-\tilde{\Sigma})v|}{\|v\|_{2}^{2}}$
$=\frac{\text{err}(\Sigma,\tilde{\Sigma})}{\|D\|_{\infty}}$
@@
Similarly, since $D_{ii}\geq(1-\epsilon)\|D\|_{\infty}$ for all $i\in[d]$, we can bound
$\text{err}_{t}(D^{-1/2}\Sigma D^{-1/2})\leq\frac{\text{err}_{t}(\Sigma)}{(1-\epsilon)\cdot\|D\|_{\infty}}.$
Putting these two results together with the bound from (12) gives the lemma.
C.2 Assumptions of Lemma 2
  1. 1.
    The $(t+1)$-th eigenvalue of $D^{-1/2}\Sigma D^{-1/2}$ is greater than or equal to $2$. In particular, letting $D^{-1/2}\Sigma D^{-1/2}=Q(I+\Lambda)Q^{\top}$ be the orthogonal eigendecomposition of the symmetrization, we assume $1+\Lambda_{t+1}\geq 2$.
  2. 2.
    The diagonal of $\Sigma$ has the property that $\min_{i\in[d]}\Sigma_{ii}\geq(1-\epsilon)\max_{j\in[d]}\Sigma_{jj}$ for some $\epsilon\in(0,1]$.
First assumption.
Notice that since $D^{-1/2}\Sigma D^{-1/2}=I+D^{-1/2}RD^{-1/2}$ is a symmetric PSD matrix, we have that $\text{tr}(D^{-1/2}\Sigma D^{-1/2})=\sum_{i=1}^{d}1+\Lambda_{i}=d$. Hence by definition there must exist some $t$ for which $1+\Lambda_{t+1}\geq 1$. While in the worse case we cannot hope for the existence of an eigenvalue larger than this, in practice, including for the datasets we consider in this work, it can be shown that in fact the eigenvalues of $D^{-1/2}\Sigma D^{-1/2}$ are larger than $2$ across shards and datasets – see Figure 5.
@@ C.2 Assumptions of Lemma 2
Figure 5: Histogram of the $(t+1)$-th eigenvalue for each dataset. For each dataset, we pick the partitioning and $t$ which we use in the plots from Figure 6. Plots show that almost all shards for all datasets have $(t+1)$-th eigenvalue bounded away from $2$, except for some shards for GloVe and one for Nq-Ada2.
@@
Since $\Sigma=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)(x_{i}-\mu)^{\top}$ for the $n$ points $x_{1},\dots,x_{n}\in\operatorname*{\mathbb{R}}^{d}$ and mean $\mu=\sum_{i=1}^{n}x_{i}/n$ in the shard, it is in fact positive semi-definite. In particular the entries of the diagonal are non-negative. While in the worst case, the diagonal can have arbitrarily large entries compared to its smallest entries, in practice this is rarely the case. While we do not explore how to remove this assumption, there are several mechanisms to do so in practice such as applying a random rotations or pseudo-random rotations Ailon and Chazelle [2009], Woodruff [2014], Ailon and Liberty [2013] to the datapoints $x_{1},\dots,x_{n}$ in each shard before using Algorithm 1. It is well known (e.g. Lemma 1 in Ailon and Chazelle [2009]) that after applying such transforms that the coordinates of the vectors are “roughly equal,” thereby ensuring that the diagonal of the covariance has entries of comparable magnitude. We leave the exploration of removing this assumption to future work.
Appendix D Additional experimental results
This section presents our full experimental results for completeness.
D.1 Clustering by Spherical KMeans
@@
Figure 6: Top-$k$ recall versus the number of probed data points. Datasets are partitioned with Spherical KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
(d) Text2Image
(e) MsMarco-MiniLM
(f) DeepImage
Figure 6: Top-$k$ recall versus the number of probed data points. Datasets are partitioned with Spherical KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
D.2 Clustering by Standard KMeans
(a) Nq-Ada2
(b) Music
(c) GloVe
Figure 7: Top-$k$ recall versus the number of probed data points. Datasets are partitioned with Standard KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
(d) Text2Image
(e) MsMarco-MiniLM
(f) DeepImage
Figure 7: Top-$k$ recall versus the number of probed data points. Datasets are partitioned with Standard KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
D.3 Clustering by GMM
We note that, due to the dimensionality of the Nq-Ada2 dataset, we were unable to complete GMM clustering on this particular dataset.
(a) Music
(b) GloVe
Figure 8: Top-$k$ recall versus the number of probed data points. Datasets are partitioned with Gaussian Mixture Model. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
(c) Text2Image
(d) MsMarco-MiniLM
(e) DeepImage
Figure 8: Top-$k$ recall versus the number of probed data points. Datasets are partitioned with Gaussian Mixture Model. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
Appendix E Effect of parameters
Recall that the final Optimist algorithm takes two configurable parameters: $t$, the rank of the covariance sketch in Equation (7); and $\delta$, the degree of optimism in Problem 1. In this section, we examine the effect of these parameters on the performance of Optimist.
In the same figure, we have also included two configurations of SubPartition: one with just $2$ sub-partitions, denoted by $\textsc{SubPartition}(0)$, and another with $t+2$ sub-partitions, $\textsc{SubPartition}(t)$. These help put the performance of Optimist with various ranks in perspective. In particular, we give the SubPartition baseline the same amount of information and contrast its recall with Optimist.

### v3

@@ 2.5. Proofs and justification of assumptions
Let $\Sigma=D+R$ be the decomposition of the PSD covariance matrix $\Sigma$ into its diagonal $D$ and residual $R=\Sigma-D$. Let $Q\Lambda Q^{\top}$ be the orthogonal eigendecomposition of $R_{\circ}=D^{-1/2}RD^{-1/2}$. Recall that we define $\Sigma_{t}^{\textsc{ms}}$, for some $1\leq t\leq d$, as
$\Sigma_{t}^{\textsc{ms}}:=D+D^{{1}/{2}}[Q]_{t}[\Lambda]_{t}[Q]_{t}^{\top}D^{{1}/{2}}$
We start by proving Lemma 3. We then proceed to justify the assumptions of the lemma in the subsequent section.
2.5.1. Proof of Lemma 3
@@
We state a few technical results that will simplify the proof of the lemma.
@@
Assuming the eigenvalues $\Lambda$ of $R_{\circ}$ are sorted in non-increasing order, we have that $I+R_{\circ}=Q(I+\Lambda)Q^{\top}$. In words, the eigenvectors of $I+R_{\circ}$ are the same as $R_{\circ}$, and the $i$-th eigenvalue is $\Lambda_{i}+1$. As a corollary, since $I+R_{\circ}$ is PSD, we have that $\Lambda_{i}\geq-1$ for all $i\in[d]$.
@@
Proof of Lemma 3.
First, note that, by the definition of $\text{err}(\cdot,\cdot)$ from Equation (7), we can write:
$\text{err}(D^{-\frac{1}{2}}\Sigma D^{-\frac{1}{2}},D^{-\frac{1}{2}}\Sigma_{t}^{\textsc{ms}}D^{-\frac{1}{2}})$ $=\sup_{v\in\operatorname*{\mathbb{R}}^{d}}\frac{|v^{\top}D^{-1/2}(\Sigma-\Sigma\Sigma_{t}^{\textsc{ms}})D^{-1/2}v|}{\|v\|_{2}^{2}}.$
Since $D$ has strictly positive entries on the diagonal, we can do a change of variables, setting $u=D^{1/2}v$. Denoting $\max_{i\in[d]}D_{ii}$ by $\|D\|_{\infty}$, this gives us:
$\text{err}(D^{-\frac{1}{2}}$ $\Sigma D^{-\frac{1}{2}},D^{-\frac{1}{2}}\Sigma_{t}^{\textsc{ms}}D^{-\frac{1}{2}})=\sup_{v\in\operatorname*{\mathbb{R}}^{d}}\frac{|v^{\top}(\Sigma-\Sigma_{t}^{\textsc{ms}})v|}{\|D^{1/2}v\|_{2}^{2}}$
$\geq\frac{1}{\|D\|_{\infty}}\cdot\sup_{v\in\operatorname*{\mathbb{R}}^{d}}\frac{|v^{\top}(\Sigma-\Sigma_{t}^{\textsc{ms}})v|}{\|v\|_{2}^{2}}=\frac{\text{err}(\Sigma,\Sigma_{t}^{\textsc{ms}})}{\|D\|_{\infty}}$
@@
Next, consider the following:
$\text{err}(D^{-\frac{1}{2}}$ $\Sigma D^{-\frac{1}{2}},D^{-\frac{1}{2}}\Sigma_{t}^{\textsc{ms}}D^{-\frac{1}{2}})=\sup_{\lVert v\rVert_{2}=1}\big{\lvert}v^{\top}D^{-\frac{1}{2}}(\Sigma-\Sigma_{t}^{\textsc{ms}})D^{-\frac{1}{2}}v\big{\rvert}$
(9) $=\sup_{\lVert v\rVert_{2}=1}\big{\lvert}v^{\top}(R_{\circ}-[Q]_{t}[\Lambda]_{t}[Q]_{t}^{\top})v\big{\rvert}$
(10) $=\sup_{\lVert v\rVert_{2}=1}\big{\lvert}\sum_{i=t+1}^{d}\Lambda_{i}\cdot\langle Q_{i},v\rangle^{2}\big{\rvert}$
(11) $=\max_{l\in[t+1,d]}\lvert\Lambda_{l}\rvert.$
where the second equality follows by the definition of $\Sigma_{t}^{\textsc{ms}}$ and the third by Fact 1.
Putting our arguments thus far together, we have established that:
(12) $\frac{\text{err}(\Sigma,\Sigma_{t}^{\textsc{ms}})}{\|D\|_{\infty}}\leq\max_{l\in[t+1,d]}\lvert\Lambda_{l}\rvert.$
Using this result, and noting that, $1+\Lambda_{t+1}\geq 1$ by assumption, so that $\Lambda_{t+1}\geq 0$; and $\Lambda_{l}\geq-1$ for all $l\in[d]$ by Fact 2, we can derive the following:
(13) $\frac{\text{err}(\Sigma,\Sigma_{t}^{\textsc{ms}})}{\|D\|_{\infty}}$ $\leq\max_{l\in[t+1,d]}|\Lambda_{l}|$
(14) $\leq 1+\Lambda_{t+1}$
(15) $\leq\max_{l\in[t+1,d]}|1+\Lambda_{l}|$
(16) $=\text{err}(\underbrace{I+R_{\circ}}_{D^{-1/2}\Sigma D^{-1/2}},[Q]_{t}[I+\Lambda]_{t}[Q]_{t}^{\top}).$
All that is left to do is to bound the right-hand side of Equation (16). We do so as follows:
(17) $\text{err}(D^{-1/2}$ $\Sigma D^{-1/2},[Q]_{t}[I+\Lambda]_{t}[Q]_{t}^{\top})$
(18) $\leq\text{err}(D^{-1/2}\Sigma D^{-1/2},D^{-1/2}\Sigma_{t}^{\textsc{lr}}D^{-1/2})$
(19) $=\sup_{\lVert v\rVert_{2}=1}\big{\lvert}v^{\top}D^{-1/2}(\Sigma-\Sigma_{t}^{\textsc{lr}})D^{-1/2}v\big{\rvert}$
(20) $=\sup_{\lVert v\rVert_{2}=1}\big{\lvert}\sum_{i,j}v_{i}v_{j}(\Sigma-\Sigma_{t}^{\textsc{lr}})_{ij}/\sqrt{D_{ii}D_{jj}}\big{\rvert}$
(21) $\leq\sup_{\lVert v\rVert_{2}=1}\big{\lvert}\sum_{i,j}v_{i}v_{j}(\Sigma-\Sigma_{t}^{\textsc{lr}})_{ij}\frac{1}{(1-\epsilon)\lVert D\rVert_{\infty}}\big{\rvert}$
(22) $\leq\frac{1}{(1-\epsilon)\lVert D\rVert_{\infty}}\sup_{\lVert v\rVert_{2}=1}\big{\lvert}v^{\top}(\Sigma-\Sigma_{t}^{\textsc{lr}})v\big{\rvert}$
(23) $=\frac{1}{(1-\epsilon)\lVert D\rVert_{\infty}}\text{err}(\Sigma,\Sigma_{t}^{\textsc{lr}}),$
where the first inequality is because the left-hand side is the optimal error by the Eckhart-Young-Mirsky Theorem, and the second inequality is due to the fact that $D_{ii}\geq(1-\epsilon)\|D\|_{\infty}$.
Putting everything together, we have shown that:
$\frac{\text{err}(\Sigma,\Sigma_{t}^{\textsc{ms}})}{\|D\|_{\infty}}\leq\frac{1}{(1-\epsilon)\lVert D\rVert_{\infty}}\text{err}(\Sigma,\Sigma_{t}^{\textsc{lr}}),$
as desired. ∎
@@
Figure 2. Histogram of the $(t+1)$-th eigenvalue. For each dataset, we pick the partitioning and $t$ from Table 1. Plots show that almost all shards for all datasets have $(t+1)$-th eigenvalue bounded away from $1$, except for a few shards for Nq-Ada2.
2.5.2. Assumptions of Lemma 3
  1. (1)
    The $(t+1)$-th eigenvalue of $D^{-1/2}\Sigma D^{-1/2}$ is greater than or equal to $1$. In particular, letting $D^{-1/2}\Sigma D^{-1/2}=Q(I+\Lambda)Q^{\top}$ be the orthogonal eigendecomposition of the symmetrization, we assume $1+\Lambda_{t+1}\geq 1$.
  2. (2)
    The diagonal of $\Sigma$ has the property that $\min_{i\in[d]}\Sigma_{ii}\geq(1-\epsilon)\max_{i\in[d]}\Sigma_{ii}$ for some $\epsilon\in(0,1]$.
First assumption.
Notice that $D^{-1/2}\Sigma D^{-1/2}=I+D^{-1/2}RD^{-1/2}$ is symmetric PSD, so we have that $\text{tr}(D^{-1/2}\Sigma D^{-1/2})=\sum_{i=1}^{d}1+\Lambda_{i}=d$. Hence by definition there must exist some $t$ for which $1+\Lambda_{t+1}\geq 1$. While in the worse case we cannot hope for the existence of an eigenvalue larger than this, in practice, including for the datasets we consider in this work, it can be shown that in fact the eigenvalues of $D^{-1/2}\Sigma D^{-1/2}$ are larger than $1$ across shards and datasets–see Figure 2.
@@ 2.5.2. Assumptions of Lemma 3
While in the worst case, the diagonal of $\Sigma$ can have arbitrarily large entries compared to its smallest entries, in practice this is rarely the case. While we do not explore how to remove this assumption, there are several mechanisms to do so in practice such as applying random rotations or pseudo-random rotations (Ailon and Chazelle, 2009; Woodruff, 2014; Ailon and Liberty, 2013) to the data points in each shard before using Algorithm 1. It is well known (e.g. Lemma 1 in (Ailon and Chazelle, 2009)) that after applying such transforms, the coordinates of the vectors are “roughly equal,” thereby ensuring that the diagonal of the covariance has entries of comparable magnitude. We leave the exploration of removing this assumption to future work.
3. Experimental evaluation
We put our arguments to the test and evaluate Optimist.
3.1. Setup
Datasets: We use the following benchmark datasets in this work:
  • Text2Image: A cross-modal dataset, where data and query points may have different distributions in a shared space (Simhadri et al., 2022). We use a subset consisting of $10$ million $200$-dimensional data points along with a subset of $10{,}000$ test queries;
  • Music: $1$ million $100$-dimensional points (Morozov and Babenko, 2018) with $1{,}000$ queries;
  • DeepImage: Subset of $10$ million $96$-dimensional points from the billion deep image features dataset (Yandex and Lempitsky, 2016) with $10{,}000$ queries;
  • GloVe: 1.21.21{.}21.2 million, $200$-dimensional word embeddings trained on tweets (Pennington et al., 2014) with $10{,}000$ test queries;
  • MsMarco-MiniLM: Ms Marco Passage Retrieval v1 (Nguyen et al., 2016) is a question-answering dataset consisting of 8.88.88{.}88.8 million short passages in English. We use the “dev” set of queries for retrieval, made up of $6{,}980$ questions. We embed individual passages and queries using the all-MiniLM-L6-v2 model [^6] Checkpoint at https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. to form a $384$-dimensional vector collection; and,
  • Nq-Ada2: 2.72.72{.}72.7 million, $1{,}536$-dimensional embeddings of the Natural Questions dataset (Kwiatkowski et al., 2019) with the Ada-002 model. [^7] https://openai.com/index/new-and-improved-embedding-model/
Table 1. Dataset statistics (size $m$ and number of dimensions $d$), along with the number of partitions ($C$) and Optimist’s default rank configuration ($t$) in our main experiments.
Dataset $m$ $d$ $C$ $t$
Text2Image $10$M $200$ $3{,}000$ $4$
Music $1$M $001$,02410241{,}0241 , 024 $2$
DeepImage $10$M $96$ $3{,}000$ $2$
GloVe $1.2$M $200$ $1{,}024$ $4$
MsMarco-MiniLM $8.8$M $843$,00030003{,}0003 , 000 $8$
Nq-Ada2 $2.7$M $1{,}536$ $1{,}600$ $30$
Clustering: For our main results, we partition datasets with Spherical KMeans (Dhillon and Modha, 2001). We include in Appendix A results from similar experiments but where the clustering algorithm is standard KMeans and Gaussian Mixture Model (GMM). We cluster each dataset into $\sqrt{m}$ shards, where $m$ is the number of data points in the dataset; this is summarized in Table 1.
Evaluation: Once a dataset has been partitioned, we fix the partitioning and evaluate all routers on it to facilitate a fair comparison. We evaluate each router $\tau$ as follows. For each test query, we identify the set of shards to probe using $\tau$. We then perform exact search over the selected shards, obtain the top-$k$ points, and compute recall with respect to the ground-truth top-$k$ set. Because the only source of error is the router’s inaccuracy, the measured recall gauges the effectiveness of $\tau$.
Routers: We evaluate the following routers in our experiments:
  • Mean and NormalizedMean: Defined in Equations (2) and (3);
  • SubPartition $(t)$: Recall that Optimist stores $t+2$ vectors per partition, where $t$ is the rank in Algorithm 1. We ask if simply partitioning each shard independently into $t+2$ sub-partitions, and recording the sub-partitions’ centroids as the representatives of that shard attains the same routing accuracy as Optimist. At query time, we take the maximum inner product of the query with a shard’s representatives as its score, and sort shards according to this score; and,
  • Optimist $(t,\delta)$: The Optimist router given in Algorithm 1. The parameter $t$ determines the rank of the sketch of the covariance matrix, and $\delta$ determines the degree of optimism. We set $\delta=0.8$ and $t$ to a maximum of $2\%$ of $d$, as summarized in Table 1, but study their effect later. Throughout our discussion, if we do not specify the parameters, it should be understood that Optimist is configured with default $\delta$ and $t$. We write Optimist $(t=d,\cdot)$ to indicate that the full covariance matrix is used.
Code: We have implemented all baseline and proposed routers in the Rust programming language. We have open-sourced [^8] Available at https://github.com/Artificial-Memory-Lab/optimist-router our code along with experimental configuration to facilitate reproducibility.
@@ 3.1. Setup
Figure 3. Top-$k$ recall vs. volume of probed data points, on Nq-Ada2, Music, and GloVe. Partitioning is by Spherical KMeans. Scann has parameter $T$, SubPartition $t$ (leading to $t+2$ sub-partitions per shard), and Optimist rank $t$ and degree of optimism $\delta$.
(a) Text2Image
(b) MsMarco-MiniLM
(c) DeepImage
Figure 4. Top-$k$ recall vs. volume of probed data points, on Text2Image, MsMarco-MiniLM, and DeepImage. Setup is the same as Figure 3.
3.2. Main results
Figures 3 and 4 plot top-$k$ recall for $k\in\{1,10,100\}$, versus the percentage of data points examined for all datasets. Partitioning is by Spherical KMeans; for standard KMeans and GMM see Figures 9 and 10.
We summarize key observations. First, among baselines, NormalizedMean generally outperforms Mean and Scann, except Music where Mean reaches a higher recall. Second, with very few exceptions, Optimist with the full covariance ($t=d$) does at least as well as NormalizedMean, and often outperforms it significantly. Interestingly, the gap between baselines and Optimist widens as retrieval depth ($k$) increases; a phenomenon that is not surprising.
Finally, while Optimist with a rank-$t$ Masked Sketch shows some degradation in comparison with no sketching (i.e., $t=d$), it still achieves a higher recall than baselines for larger $k$. When $k$ is smaller, SubPartition $(t)$ becomes a strong competitor.
Optimist shines when data points have varying norms. On Music, at $95\%$ top-$100$ recall, Optimist needs to probe $54\%$ fewer data points than NormalizedMean; on average Optimist probes $66{,}6$ 666 data points to reach $95\%$ top-$100$ recall whereas NormalizedMean examines $14{,}463$ points.
Table 2 presents savings on all datasets. On DeepImage, Optimist scans $9\%$ more data points than NormalizedMean to reach $90\%$ recall ($55{,}000$ points vs $60{,}000$), and $10\%$ more to reach $95\%$ ($95{,}000$ vs $105{,}000$). We suspect this is an artifact of the dataset’s construction: each point is represented by the top $96$ principal components of its original features, leading to unusual partition statistics.
Table 2. Comparison of Optimist and NormalizedMean by relative amount of data points saved during search to achieve a fixed recall. A saving of $x\%$ means that Optimist searches $x\%$ fewer points than NormalizedMean.
Recall Nq-Ada2 GloVe MsMarco-MiniLM Music Text2Image
$90\%$ $18\%$ $11\%$ $22\%$ $38\%$ $23\%$
$95\%$ $20\%$ $5.5\%$ $7.7\%$ $54\%$ $22\%$
3.2.1. Latency
As Table 2 shows, Optimist reaches $95\%$ recall by processing up to $54\%$ fewer points. We now study the impact of the improved routing accuracy on query latency, which consists of routing, fetching the chosen shards from storage, and scoring.
Concretely, we isolate routing and measure the elapsed time between when a router receives a query and when it returns a list of shards. This is the “routing” latency. We then measure the time it takes to fetch the chosen shards from storage; the “fetch” latency. Finally, we measure the latency of the scoring phase. We note that, for latency experiments only, we compress data points and compute inner products using PQ with $4$-dimensional codebooks.
We run experiments on an AWS c5.xlarge machine with $4$ vCPUs, $8$GB of main memory. Baseline bandwidth of the SSD attached to this machine is $1{,}150$Mbps. [^9] See official benchmark from Amazon at https://docs.aws.amazon.com/AWSEC2/latest/UserGuide/ebs-optimized.html Finally, using $4$ threads, transferring $4$MB of data from blob storage (hosted on Amazon S3) to main memory has P50 latency of $45$ milliseconds (ms). [^10] See https://github.com/dvassallo/s3-benchmark for an independent benchmark.
Refer to caption Refer to caption Refer to caption
(a) $1$ thread
(b) $2$ threads
(c) $4$ threads
Figure 5. Mean per-query latency (milliseconds) to reach $95\%$ recall, using different number of threads (columns), when data is stored on SSD (top row) and blob storage (bottom row). For each dataset, we plot the latency breakdown for NormalizedMean (top bar) and Optimist (bottom bar), and report relative gains (negative values indicate improvements, whereas positive values indicate losses) achieved by Optimist with respect to NormalizedMean. All shards are stored as PQ-compressed blobs using $4$-dimensional codebooks—a common configuration.
Figure 5 reports the routing, fetch, and scoring latency for Optimist and NormalizedMean routers. As the figure shows, Optimist’s gains are more pronounced when shards are stored on blob storage, but even on SSD the gains are substantial.
We note that, hardware plays an outsize role in latency improvements. Depending on the specifications, we may have a slower or faster interface with blob or SSD storage; a more limited bandwidth would boost Optimist’s standing in latency comparisons. On the other hand, if storage bandwidth is abundant and number of threads very limited, Optimist’s advantage would become less significant.
There are other factors that may affect the overall latency comparison. These include: target recall level; type of compression; and algorithm used for scoring. In the above experiments, we put Optimist at a disadvantage by incorporating heavy compression and fast score computation; Optimist fares even better if data remains uncompressed (as Optimist needs to transfer a smaller amount of data) or scoring is slower (as Optimist must score fewer points).
3.2.2. Router size
As we noted, a router such as NormalizedMean that uses a point estimate to score shards needs to store only a single $d$-dimensional vector per shard. Optimist, on the other hand, requires $(t+2)$ vectors in $\mathbb{R}^{d}$ per shard to make its routing decision. Here, we present concrete numbers for completeness, and compare the size of the two routers on the datasets used in our experiments.
On most datasets, the difference in size between NormalizedMean and Optimist is negligible. On GloVe, for example, NormalizedMean takes up $0.8$MB of space while Optimist $(t=4)$ consumes $4.7$MB in total. Statistics on other datasets are as follows: on MsMarco-MiniLM it is $4.4$MB for NormalizedMean versus $44$MB for Optimist $(t=8)$; on Music, $0.4$MB versus $1.6$MB ($t=2$); on Text2Image, $2.3$MB versus $13.7$MB ($t=4$); on DeepImage, $1.1$MB versus $4.4$MB ($t=2$). The difference is larger, but still mild, on Nq-Ada2: $9.4$MB versus $300$MB ($t=30$).
3.3. Effect of parameters on Optimist